OpenStudy (anonymous):

N x 4/5 = 3/4 can anyone tell me what n ='s

OpenStudy (anonymous):

\[ \frac{ 4 }{ 5 }N = \frac{ 3 }{ 4 }\] How do you isolate the N?

OpenStudy (anonymous):

N x 4/5 = 3/4